# Scale Equivariant Graph Metanetworks

### v1

@@ Abstract
This paper pertains to an emerging machine learning paradigm: learning higher-order functions, i.e. functions whose inputs are functions themselves, particularly when these inputs are Neural Networks (NNs). With the growing interest in architectures that process NNs, a recurring design principle has permeated the field: adhering to the permutation symmetries arising from the connectionist structure of NNs. However, are these the sole symmetries present in NN parameterizations? Zooming into most practical activation functions (e.g. sine, ReLU, tanh) answers this question negatively and gives rise to intriguing new symmetries, which we collectively refer to as scaling symmetries, that is, non-zero scalar multiplications and divisions of weights and biases. In this work, we propose Scale Equivariant Graph MetaNetworks - ScaleGMNs, a framework that adapts the Graph Metanetwork (message-passing) paradigm by incorporating scaling symmetries and thus rendering neuron and edge representations equivariant to valid scalings. We introduce novel building blocks, of independent technical interest, that allow for equivariance or invariance with respect to individual scalar multipliers or their product and use them in all components of ScaleGMN. Furthermore, we prove that, under certain expressivity conditions, ScaleGMN can simulate the forward and backward pass of any input feedforward neural network. Experimental results demonstrate that our method advances the state-of-the-art performance for several datasets and activation functions, highlighting the power of scaling symmetries as an inductive bias for NN processing.
@@ 1 Introduction
  • Our method is evaluated on 3 activations: ReLU (positive scale), tanh (sign) and sine (sign - we first characterise this using a technique from [25]) on several datasets and two tasks (INR classification & generalisation prediction), demonstrating superior performance against common metanetwork baselines.
@@ 2 Related work
In contrast to the above methods, our work follows a recent stream of research focusing on equivariant metanetworks. Navon et al. [52] and Zhou et al. [81] first characterised all linear equivariant layers to permutation symmetries of MLPs and combined them with non-linearities, while in the latter this was extended to CNNs. These approaches derive intricate weight-sharing patterns but are non-local222 They can be perceived as the metanetwork analogue of Invariant Graph Nets [48] for conventional graphs. and cannot process varying architectures. In follow-up works, Zhou et al. [82] constructed equivariant attention layers and in [83] the above characterisation was generalised to arbitrary input architectures and layers (e.g. RNNs and transformers) introducing an algorithm for automatic linear layer (with weight-sharing) construction. A different route was taken in the very recent works of Kofinas et al. [33] and Lim et al. [44], where input NNs are treated as a special class of graphs and are naturally processed with Graph Neural Networks (GNNs), with appropriate symmetry breaking wherever necessary. This perspective has been adopted several times in the deep learning literature [9, 23, 45, 73, 76] and in a few examples, GNNs were applied on the graph representation for, e.g. neural architecture search ([78, 69, 31]), albeit without mapping parameters to the graph edges.
@@ 5 Scale Equivariant Graph MetaNetworks
$\mathsf{ScaleInv}^{k}(\mathbf{X})=\rho^{k}\left(\mathbf{x}_{1},\dots,\mathbf{x}_{n}\right),$ (Scale Inv. Net)
@@
where $\rho^{k}:\prod_{i=1}^{n}\mathcal{X}_{i}\to\mathbb{R}^{\sum_{i=1}^{n}d_{i}^{k}}$ universal approximators (e.g. MLPs), $\mathbf{\Gamma}^{k}_{i}:\mathcal{X}_{i}\to\mathbb{R}^{d_{i}^{k}}$ linear transforms (for each of the $k$ invariant/equivariant layers resp). In practice, we observed experimentally that a single layer $K=1$ was sufficient.
@@ 6 Experiments
StatNN [70] $0.9140\pm 0.001$ $0.8463\pm 0.004$ $0.9140\pm 0.000$ $0.8440\pm 0.001$
NFNNP [81] $0.9190\pm 0.010$ $0.8586\pm 0.003$ $0.9251\pm 0.001$ $0.8580\pm 0.004$
NFNHNP [81] 0.9270 $\pm$ 0.001 0.8636 $\pm$ 0.002 $0.9339\pm 0.000$ $0.8586\pm 0.004$
NG-GNN ([33]) $0.9010\pm 0.060$ $0.8549\pm 0.002$ 0.9340 $\pm$ 0.001 0.8620 $\pm$ 0.003
ScaleGMN (Ours) 0.9276 $\pm$ 0.002 0.8689 $\pm$ 0.003 0.9418$\pm$ 0.005 0.8736 $\pm$ 0.003
ScaleGMN-B (Ours) 0.9282 $\pm$ 0.003 0.8651 $\pm$ 0.001 0.9425 $\pm$ 0.004 0.8655 $\pm$ 0.004
@@
Datasets. We evaluate ScaleGMN on datasets containing NNs with three popular activation functions: sine, ReLU and tanh. The former is prevalent in INRs, which in turn are the most appropriate testbed for GMNs. We experiment with the task of (1) INR classification, i.e. classifying functions (signals) represented as INRs. Additionally, ReLU and tanh are common in neural classifiers/regressors. A popular GMN benchmark for those is (2) generalisation prediction, i.e. predicting a classifier’s test accuracy. Here, classifiers instead of FFNNs are typically CNNs (for computer vision tasks) and to this end, we extend our method to the latter in Section A.3. In both cases, we deal with invariant tasks since the prediction of a higher-order function should remain unaffected by function-preserving weight transformations We use existing datasets that have been constructed by the authors of relevant methods, are publicly released and have been repeatedly used in the literature (8 datasets in total, 4 for each task). Finally, we follow established protocols: we perform a hyperparameter search and use the best-achieved validation metric throughout training to select our final model. We report the test metric on the iteration where the best validation is achieved.
@@
Baselines. We compare with several metanetworks in the literature. DWS [52] is a non-local metanetwork that uses all linear permutation equivariant/invariant layers interleaved with non-linearities. NFNHNP [81] is mathematically equivalent to DWS, while NFNNP [81] makes stronger symmetry assumptions in favour of parameter efficiency. The two NFN variants are also designed to process CNNs contrary to DWS. NG-GNN [33] converts each input NN to a graph (similar to our work) and employs a GNN (in specific PNA [15]). Importantly, the last three methods transform the input parameters with random Fourier features while all methods perform input normalisations to improve performance and facilitate training. These tricks are in general not equivariant to scaling and were unnecessary in the ScaleGMN experiments (more details in Section A.4). Finally, on INR classification we include a naive MLP on the flattened parameter vector and Inr2Vec [47], a non-equivariant method used for INR classification. For generalization prediction, we also compare to StatNN [70], which predicts NN accuracy based on statistical features of its weights/biases.
@@
INR classification. We design a ScaleGMN with permutation & sign equivariant components (and invariant readout). Additionally, we experiment with the bidirectional version denoted as ScaleGMNB. We use the following datasets of increasing difficulty: MNIST INR, F-MNIST INR (grayscale images) and CIFAR10-INR, Aug. CIFAR10-INR.: INRs representing images from the MNIST [38], FashionMNIST [72] and CIFAR [35] datasets resp. and an augmented version of CIFAR10-INR containing 20 different INRs for each image, trained from different initialisations (often called views). The reported metric is test accurcy. Note that [81, 82] use only the augmented dataset and hence we rerun all baselines for the original version. All input NNs correspond to functions $u_{G,\bm{\theta}}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{c}$, i.e. pixel coordinate to GS/RGB value. Further implementation details can be found in Section A.4. As shown in Table 1, ScaleGMN consistently outperforms all baselines in all datasets considered, with performance improvements compared to the state-of-the-art ranging from approx. 3% (CIFAR-10) to 13% (F-MNIST). While previous methods often resort to additional engineering strategies, such as probe features and advanced architectures [33] or extra training samples [81] to boost performance, in our case this was possible using vanilla ScaleGMNs. For example, in MNIST and F-MNIST, NG-GNN achieves 94.7±0.3 and 74.2±0.4 with 64 probe features which is apprpx. 2% and 6% below our best performance. The corresponding results for NG-T [33] (transformer) were 97.3±0.2 74.8±0.9, still on par or approx. 6% below ScaleGMN. Note that all the above are orthogonal to our work and can be used to further improve performance.
@@
Predicting CNN Generalization from weights As in prior works. we consider datasets of image classifiers and measure predictive performance using Kendall’s $\tau$ [30]. We select the two datasets used in [81], namely CIFAR-10-GS and SVHN-GS originally from Small CNN Zoo [70]. These contain CNNs with ReLU or tanh activations, which exhibit scale and sign symmetries respectively. We split each dataset into two subsets each containing the same activation and evaluate all baselines. As shown in Table 2, once again ScaleGMN outperforms all the baselines in all the examined datasets. This highlights the ability of our method to be used across different activation functions and architectures. Performance improvements here are less prominent due to the hardness of the task, a phenomenon also observed in the comparisons between prior works. Note that in this case, additional symmetries arise by the softmax classifier [36], which are currently not accounted for by none of the methods.
@@ Acknowledgements
IK, GB and YP were partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. This work was partially supported by a computational resources grant from The Cyprus Institute (HPC system “Cyclone”).
@@ A.1.5 Readout
This means that we aggregate all hidden neuron representations we aggregate all scale canonicalised/symmetrised representations of hidden with a permutation invariant aggregator (Deepsets [77]), concatenate the result with all i/o neuron representations (that cannot be permuted or scaled) and transform the result with a universal function approximator. A simpler alternative is to only concatenate the output neuron representations (or also the input for the bidirectional version) since after sufficient message-passing iterations they are expected to have collected information from the entire graph. We include and evaluate both choices in our hyperparameter grid.
@@ A.4.1 INR Classification
Hyperparameters: The GNN used for ScaleGMN is a traditional Message Passing Neural Network [24], with carefully designed update and message functions that ensure sign equivariance. We optimise the following hyperparameters: batch size in $\{64,128,256\}$, hidden dimension for node/edge features in ${64,128,256}$. We also search learning rates in $\{0.001,0.0005,0.0001\}$, weight decay in $\{0.01,0.001\}$ and dropout in $\{0,0.1,0.2\}$. Moreover, we experiment with using only vertex positional encodings or also employing edge positional encodings. We apply layer normalization within each MLP and use skip connections between each GNN layer. The last two steps were proven valuable to stabilise training. Finally, for each MLP within the architecture we use SiLU activation function, one hidden layer and no activation function after the last layer. All the experiments were conducted on NVIDIA GeForce RTX 4090.
@@
Results: As shown in table 1, our method is able to push the state-of-the-art in all the datasets considered. Notably, many additional techniques that were employed by previous works in order to achieve good results were not required in our method. Specifically, probe features when inserted into the node features as in [33] were proven to boost the performance much higher, and could potentially enhance the performance of ScaleGMN too. In the reported results we compare with variants of [33] that do not use probe features and rely solely on the parameters of the input network, although our method outperforms even the probe features variants. Moreover, [81] and [33] also apply random Fourier features during the initialization of the weight and bias features, which in our experiments showcased better performance when used in a traditional MPNN framework, but were not employed for our method. In our framework, we also do not include any normalization of the input as conducted by previous works, where the normalization is applied either on parameter level ([81], [52]) or on layer level [33], computing the mean and standard deviation. Although this step typically leads to better performance, it could not be applied in our scale and permutation equivariant framework. Nevertheless, even without all the above techniques, ScaleGMN achieves remarkable results. Finally, opting for a Transformer, as in [33], instead of a GNN layer is also an orthogonal to our work addition that could enhance the performance of ScaleGMN.

### v2

@@ Abstract
This paper pertains to an emerging machine learning paradigm: learning higher-order functions, i.e. functions whose inputs are functions themselves, particularly when these inputs are Neural Networks (NNs). With the growing interest in architectures that process NNs, a recurring design principle has permeated the field: adhering to the permutation symmetries arising from the connectionist structure of NNs. However, are these the sole symmetries present in NN parameterizations? Zooming into most practical activation functions (e.g. sine, ReLU, tanh) answers this question negatively and gives rise to intriguing new symmetries, which we collectively refer to as scaling symmetries, that is, non-zero scalar multiplications and divisions of weights and biases. In this work, we propose Scale Equivariant Graph MetaNetworks - ScaleGMNs, a framework that adapts the Graph Metanetwork (message-passing) paradigm by incorporating scaling symmetries and thus rendering neuron and edge representations equivariant to valid scalings. We introduce novel building blocks, of independent technical interest, that allow for equivariance or invariance with respect to individual scalar multipliers or their product and use them in all components of ScaleGMN. Furthermore, we prove that, under certain expressivity conditions, ScaleGMN can simulate the forward and backward pass of any input feedforward neural network. Experimental results demonstrate that our method advances the state-of-the-art performance for several datasets and activation functions, highlighting the power of scaling symmetries as an inductive bias for NN processing. The source code is publicly available at https://github.com/jkalogero/scalegmn.
@@ 1 Introduction
  • Our method is evaluated on 3 activations: ReLU (positive scale), tanh (sign) and sine (sign - we first characterise this using a technique from [25]) on several datasets and three tasks (INR classification/editing & generalisation prediction), demonstrating superior performance against common metanetwork baselines.
@@ 2 Related work
In contrast to the above methods, our work follows a recent stream of research focusing on equivariant metanetworks. Navon et al. [54] and Zhou et al. [85] first characterised all linear equivariant layers to permutation symmetries of MLPs and combined them with non-linearities, while in the latter this was extended to CNNs. These approaches derive intricate weight-sharing patterns but are non-local222 They can be perceived as the metanetwork analogue of Invariant Graph Nets [50] for conventional graphs. and cannot process varying architectures. In follow-up works, Zhou et al. [86] constructed equivariant attention layers and in [87] the above characterisation was generalised to arbitrary input architectures and layers (e.g. RNNs and transformers) introducing an algorithm for automatic linear layer (with weight-sharing) construction. A different route was taken in the very recent works of Kofinas et al. [33] and Lim et al. [44], where input NNs are treated as a special class of graphs and are naturally processed with Graph Neural Networks (GNNs), with appropriate symmetry breaking wherever necessary. This perspective has been adopted several times in the deep learning literature [9, 23, 45, 77, 80] and in a few examples, GNNs were applied on the graph representation for, e.g. neural architecture search ([82, 72, 31]), albeit without mapping parameters to the graph edges. In parallel to our work, [73] extended the framework of [85] to incorporate scaling and sign-flipping symmetries. They construct non-local equivariant layers, resulting in an architecture with fewer trainable parameters. Nevertheless, their method suffers from limited expressive power, while their experimental results showcase limited advancements. In contrast to our method, this approach limits the choice of activation functions to those equivariant to the studied symmetries. Finally, their framework cannot be extended to other activation functions, as the parameter sharing must be re-designed from scratch, while it is not suitable for processing diverse architectures.
@@ 5 Scale Equivariant Graph MetaNetworks
$\mathsf{ScaleInv}^{k}(\mathbf{X})=\rho^{k}\left(\tilde{\mathbf{x}}_{1},\dots,\tilde{\mathbf{x}}_{n}\right),$ (Scale Inv. Net)
@@
where $\rho^{k}:\prod_{i=1}^{n}\mathcal{X}_{i}\to\mathbb{R}^{\sum_{i=1}^{n}d_{i}^{k}}$ universal approximators (e.g. MLPs), $\mathbf{\Gamma}^{k}_{i}:\mathcal{X}_{i}\to\mathbb{R}^{d_{i}^{k}}$ linear transforms (for each of the $k$ invariant/equivariant layers resp - in practice, we observed experimentally that a single layer $K=1$ was sufficient), and $\tilde{\mathbf{x}}_{i}$ are explained in detail in the following paragraph.
@@ 6 Experiments
CIFAR-10-GS
both act.
StatNN [74] $0.9140\pm 0.001$ $0.8463\pm 0.004$ $0.9140\pm 0.000$ $0.8440\pm 0.001$ $0.915\pm 0.002$
NFNNP [85] $0.9190\pm 0.010$ $0.8586\pm 0.003$ $0.9251\pm 0.001$ $0.8580\pm 0.004$ $0.922\pm 0.001$
NFNHNP [85] 0.9270 $\pm$ 0.001 0.8636 $\pm$ 0.002 $0.9339\pm 0.000$ $0.8586\pm 0.004$ 0.934 $\pm$ 0.001
NG-GNN [33] $0.9010\pm 0.060$ $0.8549\pm 0.002$ 0.9340 $\pm$ 0.001 0.8620 $\pm$ 0.003 $0.931\pm 0.002$
ScaleGMN (Ours) 0.9276 $\pm$ 0.002 0.8689 $\pm$ 0.003 0.9418$\pm$ 0.005 0.8736 $\pm$ 0.003 0.941 $\pm$ 0.006
ScaleGMN-B (Ours) 0.9282 $\pm$ 0.003 0.8651 $\pm$ 0.001 0.9425 $\pm$ 0.004 0.8655 $\pm$ 0.004 0.941 $\pm$ 0.000
Table 3: Dilating MNIST INRs. MSE between the reconstructed and ground-truth image.
Method MSE in $10^{-2}$
MLP $5.35\pm 0.00$
DWS [54] $2.58\pm 0.00$
NFNNP [85] $2.55\pm 0.00$
NFNHNP [85] $2.65\pm 0.01$
NG-GNN-0 [33] $2.38\pm 0.02$
NG-GNN-64 [33] 2.06 $\pm$ 0.01
ScaleGMN (Ours) 2.56 $\pm$ 0.03
ScaleGMN-B (Ours) 1.89 $\pm$ 0.00
Table 4: Ablation: Permutation equivariant models + scaling augmentations.
Method F-MNIST
CIFAR-10-GS
ReLU
DWS [54] $67.06\pm 0.29$
DWS [54] + aug. $71.42\pm 0.45$
NFNNP [85] $68.19\pm 0.28$ $0.9190\pm 0.010$
NFNNP [85] + aug. $70.34\pm 0.13$ $0.8474\pm 0.01$
NFNHNP [85] $68.94\pm 0.64$ $0.9270\pm 0.001$
NFNHNP [85] + aug. $70.24\pm 0.47$ $0.8906\pm 0.01$
NG-GNN [33] $68.0\pm 0.20$ $0.9010\pm 0.060$
NG-GNN [33] + aug. $72.01\pm 1.4$ $0.8855\pm 0.02$
ScaleGMN (Ours) $80.46\pm 0.32$ $0.9276\pm 0.002$
ScaleGMN-B (Ours) $80.78\pm 0.16$ $0.9282\pm 0.003$
@@
Datasets. We evaluate ScaleGMN on datasets containing NNs with three popular activation functions: sine, ReLU and tanh. The former is prevalent in INRs, which in turn are the most appropriate testbed for GMNs. We experiment with the tasks of (1) INR classification (invariant task), i.e. classifying functions (signals) represented as INRs and (2) INR editing (equivariant), i.e. transforming those functions. Additionally, ReLU and tanh are common in neural classifiers/regressors. A popular GMN benchmark for those is (3) generalisation prediction (invariant), i.e. predicting a classifier’s test accuracy. Here, classifiers instead of FFNNs are typically CNNs (for computer vision tasks) and to this end, we extend our method to the latter in Section A.3. We use existing datasets that have been constructed by the authors of relevant methods, are publicly released and have been repeatedly used in the literature (8 datasets in total, 4 for each task). Finally, we follow established protocols: we perform a hyperparameter search and use the best-achieved validation metric throughout training to select our final model. We report the test metric on the iteration where the best validation is achieved.
@@
Baselines. DWS [54] is a non-local metanetwork that uses all linear permutation equivariant/invariant layers interleaved with non-linearities. NFNHNP [85] is mathematically equivalent to DWS, while NFNNP [85] makes stronger symmetry assumptions in favour of parameter efficiency. The two variants are also designed to process CNNs contrary to DWS. NG-GNN [33] converts each input NN to a graph (similar to our work) and employs a GNN (in specific PNA [15]). Importantly, the last three methods transform the input parameters with random Fourier features while all methods perform input normalisations to improve performance and facilitate training. These tricks are in general not equivariant to scaling and were unnecessary in the ScaleGMN experiments (more details in Section A.4). On INR classification we include a naive MLP on the flattened parameter vector and Inr2Vec [47], a task-specific non-equivariant method. For generalization prediction, we also compare to StatNN [74], which predicts NN accuracy based on statistical features of its weights/biases.
@@
INR classification. We design a ScaleGMN with permutation & sign equivariant components (and invariant readout). Additionally, we experiment with the bidirectional version denoted as ScaleGMNB. We use the following datasets of increasing difficulty: MNIST INR, F-MNIST INR (grayscale images) and CIFAR10-INR, Aug. CIFAR10-INR.: INRs representing images from the MNIST [38], FashionMNIST [76] and CIFAR [35] datasets resp. and an augmented version of CIFAR10-INR containing 20 different INRs for each image, trained from different initialisations (often called views). The reported metric is test accuracy. Note that [85, 86] use only the augmented dataset and hence we rerun all baselines for the original version. All input NNs correspond to functions $u_{G,\bm{\theta}}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{c}$, i.e. pixel coordinate to GS/RGB value. Further implementation details can be found in Section A.4.1. As shown in Table 1, ScaleGMN consistently outperforms all baselines in all datasets considered, with performance improvements compared to the state-of-the-art ranging from approx. 3% (CIFAR-10) to 13% (F-MNIST). While previous methods often resort to additional engineering strategies, such as probe features and advanced architectures [33] or extra training samples [85] to boost performance, in our case this was possible using vanilla ScaleGMNs. For example, in MNIST and F-MNIST, NG-GNN achieves $94.7\pm 0.3$ and $74.2\pm 0.4$ with 64 probe features which is apprpx. 2% and 6% below our best performance. The corresponding results for NG-T [33] (transformer) were $97.3\pm 0.2$ and $74.8\pm 0.9$, still on par or approx. 6% below ScaleGMN. Note that all the above are orthogonal to our work and can be used to further improve performance.
@@
Predicting CNN Generalization from weights. As in prior works. we consider datasets of image classifiers and measure predictive performance using Kendall’s $\tau$ [30]. We select the two datasets used in [85], namely CIFAR-10-GS and SVHN-GS originally from Small CNN Zoo [74]. These contain CNNs with ReLU or tanh activations, which exhibit scale and sign symmetries respectively. To assess the performance of our method (1) on each activation function individually and (2) on a dataset with heterogeneous activation functions we discern distinct paths for this experiment. In the first case, we split each dataset into two subsets each containing the same activation and evaluate all baselines. As shown in Table 2, once again ScaleGMN outperforms all the baselines in all the examined datasets. This highlights the ability of our method to be used across different activation functions and architectures. Performance improvements here are less prominent due to the hardness of the task, a phenomenon also observed in the comparisons between prior works. Note that in this case, additional symmetries arise by the softmax classifier [36], which are currently not accounted for by none of the methods. We additionally evaluate our method on heterogeneous activation functions. In principle, our method does not impose limitations regarding the homogeneity of the activation functions of the input NNs - all one needs is to have a different canonicalisation/symmetrisation module for each activation. Experiments on CIFAR-10-GS show that ScaleGMN yields superior performance compared to the baselines, significantly exceeding the performance of the next-best model. Further implementation details can be found in Section A.4.2.
INR editing. Here, our goal is to transform the weights of the input NN, to modify the underlying signal to a new one. In this case, our method should be equivariant to the permutation and scaling symmetries, such that every pair of equivalent input NNs is mapped to a pair of equivalent output NNs. Hence, we employ a ScaleGMN similar to the above experiments but omit the invariant readout layer. Following [33], we evaluate our method on the MNIST dataset and train our model to dilate the encoded MNIST digits. Further implementation details can be found in Section A.4.3. As shown in Table 4, our bidirectional ScaleGMN-B achieves an MSE test loss ($10^{-2}$) equal to $1.891$, surpassing all permutation equivariant baselines. Notably, our method also outperforms the NG-GNN [33] baseline that uses 64 probe features. Additionally, our forward variant, ScaleGMN, performs on par with the previous permutation equivariant baselines with an MSE loss of $2.56$. Note that the performance gap between the forward and the bidirectional model is probably expected for equivariant tasks: here we are required to compute representations for every node of the graph, yet in the forward variant, the earlier the layer of the node, the smaller the amount of information it receives. This observation corroborates the design choices of the baselines, which utilize either bidirectional mechanisms (NG-GNN [33]) or non-local operations (NFN [85]).
Ablation study: Scaling data augmentations. We baseline our method with permutation equivariant methods trained with scaling augmentations: For every training datapoint, at each training iteration, we sample a diagonal scaling matrix for every hidden layer of the input NN and multiply it with the weights/bias matrices as per Eq. 3 (omitting the permutation matrices). We sample the elements of the matrices independently as follows: Sign symmetry: Bernoulli distribution with probability 0.5. Positive scaling: Exponential distribution where the coefficient $\lambda$ is a hyperparameter that we tune on the validation set. Observe here that designing augmentations in the latter case is a particularly challenging task since we have to sample from a continuous and unbounded distribution. Our choice of the exponential was done by consulting the norm plots where in some cases the histogram resembles an exponential Fig. 1. Nevertheless, regardless of the distribution choice we cannot guarantee that the augmentations will be sufficient to achieve (approximate) scaling equivariance, due to the lack of upper bound. We evaluate on the F-MNIST dataset for the sign symmetry and on the CIFAR-10-GS-ReLU for the positive scale. As shown in Table 4, regarding the former, augmenting the training set leads consistently to better results when compared to the original baselines. None of these methods however achieved results on par with ScaleGMN and ScaleGMN-B. On the other hand, we were unable to even surpass the original baselines regarding the latter task. This indicates that designing an effective positive scaling augmentation might be a non-trivial task.
@@ Acknowledgements
IK, GB and YP were partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program. This work was partially supported by a computational resources grant from The Cyprus Institute (HPC system “Cyclone”) as well as from an AWS credits grant provided by GRNET – National Infrastructures for Research and Technology.
@@ A.1.5 Readout
This means that we aggregate all scale canonicalised/symmetrised representations of hidden neurons with a permutation invariant aggregator (Deepsets [81]), concatenate the result with all i/o neuron representations (that cannot be permuted or scaled) and transform the result with a universal function approximator. A simpler alternative is to only concatenate the output neuron representations (or also the input for the bidirectional version) since after sufficient message-passing iterations they are expected to have collected information from the entire graph. We include and evaluate both choices in our hyperparameter grid.
@@ A.4.1 INR Classification
Hyperparameters: The GNN used for ScaleGMN is a traditional Message Passing Neural Network [24], with carefully designed update and message functions that ensure sign equivariance. We optimise the following hyperparameters: batch size in $\{64,128,256\}$, hidden dimension for node/edge features in $\{64,128,256\}$. We also search learning rates in $\{0.001,0.0005,0.0001\}$, weight decay in $\{0.01,0.001\}$, dropout in $\{0,0.1,0.2\}$ and number of GNN layers in $\{2,3,4,5\}$. Moreover, we experiment with using only vertex positional encodings or also employing edge positional encodings. We apply layer normalization within each MLP and use skip connections between each GNN layer. The last two steps were proven valuable to stabilise training. Finally, for each MLP within the architecture we use SiLU activation function, one hidden layer and no activation function after the last layer. All the experiments were conducted on NVIDIA GeForce RTX 4090.
@@
Results: As shown in table Table 1, our method is able to push the state-of-the-art in all the datasets considered. Notably, many additional techniques that were employed by previous works in order to achieve good results were not required in our method. Specifically, probe features when inserted into the node features as in [33] were proven to boost the performance much higher, and could potentially enhance the performance of ScaleGMN too. In the reported results we compare with variants of [33] that do not use probe features and rely solely on the parameters of the input network, although our method outperforms even the probe features variants. Moreover, [85] and [33] also apply random Fourier features during the initialization of the weight and bias features, which in our experiments showcased better performance when used in a traditional MPNN framework, but were not employed for our method. In our framework, we also do not include any normalization of the input as conducted by previous works, where the normalization is applied either on parameter level ([85], [54]) or on layer level [33], computing the mean and standard deviation. Although this step typically leads to better performance, it could not be applied in our scale and permutation equivariant framework. Nevertheless, even without all the above techniques, ScaleGMN achieves remarkable results. Finally, opting for a Transformer, as in [33], instead of a GNN layer is also an orthogonal to our work addition that could enhance the performance of ScaleGMN.